# Precession and higher order multipoles in binary black holes (and lack thereof)

### v1

@@ V Acknowledgements
We are grateful to Laura Nuttall for continued discussions throughout this project, and to Eleanor Hamilton for valuable comments on this manuscript. We thank Sylvia Biscoveanu, Juan Calderón Bustillo, Thomas Callister, Mark Hannam, Cameron Mills, Alex Nitz, Soumen Roy, Stefano Schmidt, Eric Thrane, John Veitch and Barak Zackay for discussions and suggestions. CH thanks the UKRI Future Leaders Fellowship for support through the grant MR/T01881X/1. SF thanks the Science and Technology Facilities Council (STFC) for support through the grant ST/N005430/1. IM acknowledges support from the Australian Research Council (ARC) Centre of Excellence for Gravitational Wave Discovery (OzGrav), project number CE230100016. We are grateful for computational resources provided by Cardiff University, LIGO Laboratory and the ICG, SEPNet and the University of Portsmouth, and supported by STFC grant ST/N000064 and National Science Foundation Grants PHY-0757058 and PHY-0823459. This research made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO is funded by the U.S. National Science Foundation. Virgo is funded by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by Polish and Hungarian institutes. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
@@ Appendix A Assessing the harmonic decomposition of GWs
As described in Section II, the harmonic decomposition of GWs allows us to explicitly isolate the precessing and higher order multipole components of observed GWs. For most of the parameter space, a GW signal can be well approximated as a sum of the leading three waveform components, $h\approx h^{\prime}+h_{\mathrm{p}}+h_{33}$ where $h_{\mathrm{p}}$ is the precessing component and $h_{33}$ corresponds to the $(\ell,|m|)=(3,3)$ higher order multipole contribution.
In Fig 5 we compare the true and approximate waveforms for a binary with masses $40\,M_{\odot}$ and $10\,M_{\odot}$, and explicitly show $h^{\prime},h_{\mathrm{p}}$ and $h_{33}$. This binary is precessing with component spin magnitudes of $\chi=0.5$, both tilt angles between the spin and the orbital angular momentum equal to $\theta=0.8$ radians, and is viewed at an inclination angle of $0.5$ radians. When generating the true waveform we use the IMRPhenomXPHM [78] phenomenological model, which includes the effects of both precession and higher order multipoles and is regularly used by the LVK. We generate $h_{\mathrm{p}}$ and $h_{33}$ using code available in the PESummary [77] and simple-pe [70] libraries.
From Fig 5 we see that $h^{\prime}$ is the dominant contribution to the waveform with SNR $19.4$. $h_{\mathrm{p}}$ and $h_{33}$ have smaller amplitudes, with SNRs $\rho_{\mathrm{p}}=3.0$ and $\rho_{33}=3.0$. The total SNR of the approximate waveform $h^{\prime}+h_{\mathrm{p}}+h_{33}$ is $19.8$, which is slightly smaller than the true waveform’s SNR of 20. This difference is attributed to additional power in higher order multipoles beyond $(\ell,|m|)=(3,3)$. We find that although the GW is written with only 3 components, the sum provides a good approximation to the true GW. We quantify this agreement with the overlap between the true waveform, $g$, and the approximate waveform, $h$, defined as,
@@
$\tilde{h}(f)$ represents the Fourier transform of $h(t)$, * implies a complex conjugate, and $S(f)$ is the power spectral density. An overlap [^3] Note the overlap does not maximise over phase and time. close to unity implies that the waveforms are very similar, while a lower overlap implies that there are significant differences between the waveforms. Assuming a typical PSD from O3, we find that the overlap between $g$ and $h$ is $0.97$ when considering only the inspiral (the harmonic decomposition is more accurate during the inspiral than during the merger and ringdown). This overlap reduces to 0.95 when including the merger and ringdown. For this case, we chose an exotic binary that is unlikely assuming the current population of binary black holes [15]. This was chosen to exaggerate the effects of precession and higher-order multipoles. We therefore expect the overlap will be closer to 1.0 for the observed population of binary black holes.

### v2

@@ V Acknowledgements
We are grateful to Laura Nuttall for continued discussions throughout this project, and to Eleanor Hamilton for valuable comments on this manuscript. We thank Sylvia Biscoveanu, Juan Calderón Bustillo, Thomas Callister, Mark Hannam, Cameron Mills, Alex Nitz, Soumen Roy, Stefano Schmidt, Eric Thrane, John Veitch and Barak Zackay for discussions and suggestions. CH thanks the UKRI Future Leaders Fellowship for support through the grant MR/T01881X/1. SF thanks the Science and Technology Facilities Council (STFC) for support through the grant ST/V005618/1. IM acknowledges support from the Australian Research Council (ARC) Centre of Excellence for Gravitational Wave Discovery (OzGrav), project number CE230100016. We are grateful for computational resources provided by Cardiff University, LIGO Laboratory and the ICG, SEPNet and the University of Portsmouth, and supported by STFC grant ST/N000064 and National Science Foundation Grants PHY-0757058 and PHY-0823459. This research made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO is funded by the U.S. National Science Foundation. Virgo is funded by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by Polish and Hungarian institutes. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
@@ Appendix A Assessing the harmonic decomposition of GWs
As described in Section II, the harmonic decomposition of GWs allows us to explicitly isolate the precessing and HoM components of observed GWs. For most of the parameter space, a GW signal can be well approximated as a sum of the leading three waveform components, $h\approx h^{\prime}+h_{\mathrm{p}}+h_{33}$ where $h_{\mathrm{p}}$ is the precessing component and $h_{33}$ corresponds to the $(\ell,|m|)=(3,3)$ HoM contribution.
In Fig 5 we compare the true and approximate waveforms for a binary with masses $40\,M_{\odot}$ and $10\,M_{\odot}$, and explicitly show $h^{\prime},h_{\mathrm{p}}$ and $h_{33}$. This binary is precessing with component spin magnitudes of $\chi=0.5$, both tilt angles between the spin and the orbital angular momentum equal to $\theta=0.8$ radians, and is viewed at an inclination angle of $0.5$ radians. When generating the true waveform we use the IMRPhenomXPHM [83] phenomenological model, which includes the effects of both precession and HoMs and is regularly used by the LVK. We generate $h_{\mathrm{p}}$ and $h_{33}$ using code available in the PESummary [82] and simple-pe [70] libraries.
From Fig 5 we see that $h^{\prime}$ is the dominant contribution to the waveform with SNR $19.4$. $h_{\mathrm{p}}$ and $h_{33}$ have smaller amplitudes, with SNRs $\rho_{\mathrm{p}}=3.0$ and $\rho_{33}=3.0$. The total SNR of the approximate waveform $h^{\prime}+h_{\mathrm{p}}+h_{33}$ is $19.8$, which is slightly smaller than the true waveform’s SNR of 20. This difference is attributed to additional power in HoMs beyond $(\ell,|m|)=(3,3)$. We find that although the GW is written with only 3 components, the sum provides a good approximation to the true GW. This implies that Eq. 1 is valid even for regions of the parameter space that are unlikely assuming the current population of binary black holes [15]. We quantify this agreement with the overlap between the true waveform, $g$, and the approximate waveform, $h$, defined as,
@@
$\tilde{h}(f)$ represents the Fourier transform of $h(t)$, * implies a complex conjugate, and $S(f)$ is the power spectral density. An overlap [^7] Note the overlap does not maximise over phase and time. close to unity implies that the waveforms are very similar, while a lower overlap implies that there are significant differences between the waveforms. Assuming a typical PSD from O3, we find that the overlap between $g$ and $h$ is $0.97$ when considering only the inspiral (the harmonic decomposition is more accurate during the inspiral than during the merger and ringdown). This overlap reduces to 0.95 when including the merger and ringdown.
Current population models show a preference for comparable mass components with half of the spin magnitudes below 0.250.25~{}0.250.25 [62, 15, 27, 63, 64, 65, 66]. Therefore, the impact of precession and HoMs will be significantly reduced compared with the system shown in Fig 5. We therefore expect the overlap to be closer to 1.0 for the observed population of binary black holes. Indeed, when randomly drawing 1,000 observable binaries from population models, assuming current detector sensitivities, we find an average overlap between the approximate and full waveforms to be $0.98$. We therefore expect Eq. 1 to be a good approximation to the true GW for the majority of binary black holes observed by the LVK.